$SPAD/src/input roots.input

Timothy Daly December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1991.

———-

— * —

)set break resume
)spool roots.output
)set message test on
)set message auto off
)clear all

———-

This will compute all the roots of x**4 + 1 = 0 — * —

--S 1 of 7
lr:=rootsOf(x**4+1,x)
--R
--R
--R   (1)  [%x0,%x0 %x1,- %x0,- %x0 %x1]
--R                                              Type: List(Expression(Integer))
--E 1

———-

The new expressions satisfy — * —

--S 2 of 7
definingPolynomial %x0
--R
--R
--R           4
--R   (2)  %x0  + 1
--R                                                    Type: Expression(Integer)
--E 2
--S 3 of 7
definingPolynomial %x1
--R
--R
--R           2
--R   (3)  %x1  + 1
--R                                                    Type: Expression(Integer)
--E 3

———-

The product of the first three is a root — * —

--S 4 of 7
lr.1 * lr.2 * lr.3
--R
--R
--R             3
--R   (4)  - %x0 %x1
--R                                                    Type: Expression(Integer)
--E 4
--S 5 of 7
%**4
--R
--R
--R   (5)  - 1
--R                                                    Type: Expression(Integer)
--E 5

———-

So is the sum — * —

--S 6 of 7
lr.1 + lr.2 + lr.3
--R
--R
--R   (6)  %x0 %x1
--R                                                    Type: Expression(Integer)
--E 6
--S 7 of 7
%**4
--R
--R
--R   (7)  - 1
--R                                                    Type: Expression(Integer)
--E 7
)spool
)lisp (bye)

———-

References^{[1] nothing}

5